# Using Confidence Scores to Improve Eyes-free Detection of Speech Recognition Errors

### v1

@@ Abstract.
Conversational systems rely heavily on speech recognition to interpret and respond to user commands and queries. Nevertheless, recognition errors may occur, which can significantly affect the performance of such systems. While visual feedback can help detect errors, it may not always be practical, especially for people who are blind or low-vision. In this study, we investigate ways to improve error detection by manipulating the audio output of the transcribed text based on the recognizer’s confidence level in its result. Our findings show that selectively slowing down the audio when the recognizer exhibited uncertainty led to a relative increase of 12% in participants’ error detection ability compared to uniformly slowing down the audio.
@@ 1. Introduction
Prior work has investigated helping users detect and correct errors by providing visual feedback (Fujiwara, 2016; Price and Sears, 2005). However, visual feedback may not be possible for individuals with visual impairments or for sighted users in situations in which they cannot visually attend to their device. Identifying errors in conversational systems without visual feedback can be challenging for a variety of reasons. Firstly, text-to-speech (TTS) audio can be hard to understand, especially when errors involve short or similar-sounding words (Burke et al., 2006). In a study with sighted users (Hong and Findlater, 2018), participants missed approximately 50% of recognition errors when the TTS audio was played at a rate of 200 words per minute (wpm). This difficulty in understanding TTS can be made even worse when the user is in a noisy environment. Finally, errors may occur infrequently, lulling users into trusting the recognizer.
In this study, we examine how users can detect speech recognition errors through audio-only feedback. Similar to the study by Hong and Findlater (Hong and Findlater, 2018), we investigate the impact of various TTS manipulations on users’ ability to detect ASR errors. Hong and Findlater found that the ability to identify errors improved when audio feedback was delivered at a speech rate of 200 wpm, or even slower at 100 wpm, compared to a higher speech rate of 300 wpm. In this work, we investigate if users’ ability to detect errors can be improved by adjusting the audio feedback using the speech recognizer’s confidence score. The confidence score indicates how certain the ASR system is about the accuracy of its result (Gillick et al., 1997). Furthermore, we investigate participants’ ability to detect errors in both common phrases where all words were in-vocabulary, and challenging phrases where at least one word was out-of-vocabulary (e.g. acronyms, proper names).
We studied the effect of confidence scores on eyes-free error detection by testing four audio annotations: default speech rate, slow speech rate, slow speech rate only for low confidence recognitions, and the inclusion of a beep tone for low confidence recognitions. Results showed that slowing down the TTS audio based on confidence score led to 85.3% accurate error detection, outperforming uniform slowing of the audio. Despite the increased audio length resulting from selective slowing based on the confidence score, participants only experienced a slight 7% increase in the time it took them to review errors compared to the default speech rate condition.
@@ 2.1. Participants
We recruited 48 participants (15 female, 32 male) aged between 21 to 68 (M=36, SD=11.5) via Amazon’s Mechanical Turk, an online crowdsourcing platform. Participants were compensated at a rate of $10 (USD) per hour. Participants completed the experiment in 27 minutes on average. All participants self-reported being native English Speakers. 67% of the participants agreed that they frequently used speech interfaces with 22% agreeing that computers had difficulty understanding their speech. Participants rated statements using a 7-point scale with one denoting strongly disagree and seven denoting strongly agree. See Appendix A for the exact questionnaires we used.
@@ 2.2. Study Design
AllSlow — The result was played at 70% of the default speaking rate, equivalent to 140 wpm.
@@
In the UncertainSlow and UncertainBeep conditions, we slowed down the TTS audio or added a beep tone when the confidence score fell below a threshold. To establish this threshold, we conducted a pilot study with 12 participants. Using the 480 utterances collected from the pilot study, we recognized the audio using Google’s speech-to-text service111https://cloud.google.com/speech-to-text and tested different thresholds measuring two metrics: 1) the true positive rate (TPR), the proportion of utterances containing one more or recognition errors that were correctly identified as having an error, and 2) the false positive rate (FPR), the proportion of utterances with no errors that were incorrectly identified as having an error. We evaluated the trade-off between the TPR and the FPR at different thresholds with a receiver operating characteristic curve. We identified an optimal threshold of 0.93 for the confidence score, which achieved a sensitivity of 0.85 and a specificity of 0.75, indicating this threshold was able to accurately detect a high percentage of errors while avoiding false positives.
@@ 2.3. Procedure
Using a web application, participants first signed a consent form and filled up a demographic questionnaire. At the beginning of the study, participants were provided instructions and completed two practice tasks to familiarize themselves with the task. The audio was played at the default speaking rate in the practice trial. At the start of each condition, we provided participants with a description of how the audio annotation worked for the current condition.
Participants recorded a sentence for each task and the recording was transcribed by Google’s speech-to-text service, then synthesized into speech with Google’s TTS service222https://cloud.google.com/text-to-speech. Speech Synthesis Markup Language (SSML) was used in the TTS request to annotate the audio response (slowing speech rate or adding a beep). Following a delay caused by the speech-to-text and TTS processing (averaging around four seconds), participants were presented with the TTS of the recognition result. Participants were not allowed to play back the TTS audio again. Participants were then asked if the reference sentence matched the audio. If they answered no, indicating a speech recognition error, they were asked to locate the incorrect or missing words, as well as any incorrect additional words that may have appeared between two words in the reference sentence (Figure 1). Participants could only detect errors after the audio finished playing, simulating a real-world scenario where users cannot interrupt the system to correct any errors they detect. At the end of the study, participants completed a final questionnaire about their experience.
@@ 3.1. Error Detection Accuracy
We calculated how often users correctly determined whether the recognition result contained any errors (i.e. by selecting yes or no after hearing the audio). As shown in Figure 2(a), the proportion of correct error detection was higher in UncertainSlow (85%) compared to AllNormal (80%), AllSlow (76%), and UncertainBeep (79%). A non-parametric ART test revealed a significant difference ( = 4.48, = 0.087, = 0.005). Post-hoc pairwise comparisons with Bonferroni correction found a significant difference between UncertainSlow and All Slow conditions ( = 0.002). In contrast to the previous study by Hong and Findlater (Hong and Findlater, 2018) that reported improved error detection with a slow speech rate, our study did not find a significant difference in error detection between AllSlow and AllNormal conditions ( = 0.94). However, our results suggest that slowing down the audio playback only when necessary might help users to better detect the presence of errors compared to uniformly slowing down the playback.
@@
In our study, participants were presented with both common and challenging phrases. The WER was significantly higher at 17% for challenging phrases compared to 12% for common phrases ( = -0.97, ¡ 0.001). This suggests that our approach of using challenging phrases to elicit more recognition errors was effective. Unfortunately, we lacked sufficient data to reliably analyze the impact of different audio annotations on participants’ ability to identify the specific locations of the errors. This was due to not every participant experiencing a sufficient number of recognition errors in each condition. Across all data, the ratio of correctly locating errors to the actual error locations was 49%, with 2% for insertions, 49% for substitutions, and 62% for deletions. Actual substitution and deletion errors were identified by aligning the reference and recognition transcripts using the Levenshtein distance algorithm (Levenshtein et al., 1966). We determined actual insertion errors by manual review. Across all conditions, the ratio of locating errors was 48% for challenging phrases and 52% for common phrases. This indicates participants missed nearly half of the errors in the transcribed text. Particularly, they struggled with identifying insertions, suggesting that detecting and correcting added words may necessitate greater attention.
@@ 3.2. Detection Time
We measured the detection time from the end of the audio playback to the participants’ response of either yes or no. Average detection times were similar: 1.98 seconds in AllNormal, 2.03 seconds in All Slow, 1.86 seconds in Uncertain Slow, and 2.26 seconds in Uncertain Beep condition (Figure 2(b)). These differences were not significant ( = 0.69, = 0.014, = 0.56). However, when considering the total time to review errors by adding the audio playback time, a significant difference was observed ( = 14.2, = 0.23, ¡ 0.001). Participants took less time in reviewing errors in AllNormal (4.52 seconds) condition compared to UncertainSlow (4.85 seconds), AllSlow (5.43), and UncertainBeep (5.42 seconds) conditions. Post-hoc pairwise comparisons revealed significant differences between AllSlow and UncertainSlow ( ¡ 0.001) but not between UncertainSlow and AllNormal conditions ( = 0.16). This indicates that selectively slowing down the audio did not significantly increase the overall review time despite the increased audio length.
@@ 4. Discussion
In our study, we used four audio annotations to assess participants’ ability to detect errors in their transcribed speech. Our result suggests that providing users with a measure of confidence in the accuracy of the transcription can be helpful for improving error detection. One limitation of our study is that we only considered sighted native English users. Blind users, for example, may have more experience listening to TTS, which could impact their ability to detect errors. Additionally, non-native users who have accents or different pronunciations may also have different experiences with speech recognition technology. Future research should explore how diverse populations interact with speech recognition technology to detect errors.
In the real world, users may be exposed to various types of noises and distractions that could affect their ability to detect errors. Moreover, users may be engaged in other tasks while using voice assistants (e.g. driving or exercising), which could also affect their ability to detect errors. Future studies could investigate how different contexts and tasks impact users’ ability to detect speech recognition errors.
To create a realistic task, we had participants record themselves speaking sentences and used a speech recognizer to transcribe them. Unlike pre-recorded audio which always contains errors, users in real-world scenarios may not always notice small errors if the recognizer is mostly accurate. However, due to our data lacking sufficient recognition errors for each participant and condition, we were unable to analyze participants’ ability to detect the specific locations of these errors. Nevertheless, the inclusion of challenging phrases was successful in significantly increasing the WER compared to common phrases. We suggest future work consider additional ways to ensure sufficient errors in each condition such as 1) a longer or multi-session study with more utterances per condition, 2) adding noise to participant’s audio to increase errors, or 3) injecting forced errors by occasionally presenting the second best recognition result.
@@ 5. Conclusion
In conclusion, our study investigated whether audio annotations based on confidence scores helped participants detect speech recognition errors. We found annotating audio based on confidence score did facilitate better detection of error presence. We believe that our results offer valuable insights that could inform the design of more effective error detection and correction systems, thereby improving the accuracy and usability of conversational systems.

### v2

@@ Abstract.
Conversational systems rely heavily on speech recognition to interpret and respond to user commands and queries. Despite progress on speech recognition accuracy, errors may still sometimes occur and can significantly affect the end-user utility of such systems. While visual feedback can help detect errors, it may not always be practical, especially for people who are blind or low-vision. In this study, we investigate ways to improve error detection by manipulating the audio output of the transcribed text based on the recognizer’s confidence level in its result. Our findings show that selectively slowing down the audio when the recognizer exhibited uncertainty led to a 12% relative increase in participants’ ability to detect errors compared to uniformly slowing the audio. It also reduced the time it took participants to listen to the recognition result and decide if there was an error by 11%.
@@ 1. Introduction
Prior work has investigated helping users detect and correct errors by providing visual feedback (Fujiwara, 2016; Price and Sears, 2005). However, visual feedback may not be possible for individuals with visual impairments, for sighted users in situations in which they cannot visually attend to their device, or when using a screenless device such as a smart speaker. Identifying errors in conversational systems without visual feedback can be challenging for a variety of reasons. Firstly, text-to-speech (TTS) audio can be hard to understand, especially when errors involve short or similar-sounding words (Burke et al., 2006). In a study with sighted users (Hong and Findlater, 2018), participants missed approximately 50% of recognition errors when the TTS audio was played at a rate of 200 words per minute (wpm). This difficulty in understanding TTS can be made even worse when the user is in a noisy environment. Finally, errors may occur infrequently, lulling users into trusting the recognizer.
In this study, we examine how users can detect speech recognition errors through audio-only feedback. Similar to the study by Hong and Findlater (Hong and Findlater, 2018), we investigate the impact of various TTS manipulations on users’ ability to detect ASR errors. Hong and Findlater found that the ability to identify errors improved when audio feedback was delivered at a speech rate of 200 wpm, or even slower at 100 wpm, compared to a higher speech rate of 300 wpm. In this work, we investigate if we can improve error detection by adjusting the audio feedback using the speech recognizer’s confidence score. The confidence score indicates how certain the ASR system is about the accuracy of its result (Gillick et al., 1997). Furthermore, we investigate participants’ ability to detect errors in both common phrases where all words were in-vocabulary, and challenging phrases where at least one word was out-of-vocabulary (e.g. acronyms, proper names).
We studied the effect of confidence scores on eyes-free error detection by testing four audio annotations: default speech rate, slow speech rate, slow speech rate only for low confidence recognitions, and the inclusion of a beep tone for low confidence recognitions. Results showed that slowing down the TTS audio based on confidence score led to 85.3% accurate error detection, outperforming uniform slowing of the audio to 70% of the default speaking rate. Despite the increased audio length resulting from selective slowing based on the confidence score, participants only experienced a slight 7% increase in the time it took them to review errors compared to the default speech rate condition.
@@ 2.1. Participants
We recruited 48 participants (15 female, 32 male) aged between 21 to 68 (mean=36, sd=11.5) via Amazon’s Mechanical Turk, an online crowdsourcing platform. We opted for an online study in order to to ensure participant safety during the COVID-19 pandemic. Participants were compensated at a rate of $10 (USD) per hour. Participants completed the experiment in 27 minutes on average. All participants self-reported being native English Speakers. 67% of the participants agreed that they frequently used speech interfaces with 22% agreeing that computers had difficulty understanding their speech. At the end of the study, participants rated various statements using a 7-point scale with one denoting strongly disagree and seven denoting strongly agree. See Appendix A for the exact questionnaires we used.
@@ 2.2. Study Design
AllSlow — The result was played at 70% of the default speaking rate, equivalent to 140 wpm. This differs from the slow condition in the study by Hong and Findlater (Hong and Findlater, 2018), where the audio was played at a slower rate of 100 wpm.
@@
In the UncertainSlow and UncertainBeep conditions, we used a confidence threshold to determine whether to slow the TTS audio or add a beep. To establish this threshold, we conducted a pilot study with 12 participants. The pilot was conducted similar to our main study but used an initial guess for the confidence threshold. Using the 480 utterances collected from the pilot study, we recognized the audio using Google’s speech-to-text service111https://cloud.google.com/speech-to-text. We tested different thresholds measuring two metrics: 1) the true positive rate (TPR), the proportion of utterances containing one more or recognition errors that were correctly identified as having an error, and 2) the false positive rate (FPR), the proportion of utterances with no errors that were incorrectly identified as having an error. We evaluated the trade-off between the TPR and the FPR at different thresholds with a receiver operating characteristic (ROC) curve. We identified an optimal threshold of 0.93 for the confidence score. Using this threshold we achieved a sensitivity of 0.85 and a specificity of 0.75, indicating we detected a high percentage of errors while avoiding too many false positives.
@@ 2.3. Procedure
Using a web application, participants first signed a consent form and filled out a demographic questionnaire. At the beginning of the study, participants were provided instructions and then completed two practice tasks to familiarize themselves with the task. The audio was played at the default speaking rate in the practice trial. At the start of each condition, we provided participants with a description of how the audio annotation worked for the current condition.
Participants recorded a sentence for each task and the recording was transcribed by Google’s speech-to-text service, then synthesized using Google’s TTS service222https://cloud.google.com/text-to-speech. Speech Synthesis Markup Language (SSML) was used in the TTS request to annotate the audio response in order to slow the speech rate or add a beep. Following a delay caused by the speech-to-text and TTS processing (averaging around four seconds), participants were presented with the audio of the recognition result. Participants were not allowed to play back the audio again. Participants were then asked if the reference sentence matched the audio. If they answered no, indicating a speech recognition error, they were asked to locate the incorrect or missing words, as well as any incorrect additional words that may have appeared between two words in the reference sentence (Figure 1). Participants could only detect errors after the audio finished playing, simulating a real-world scenario where users cannot interrupt the system to correct any errors they detect during the initial playback. At the end of the study, participants completed a final questionnaire about their experience in each condition and in the study as a whole.
@@ 3.1. Error Detection Accuracy
We calculated how often users correctly determined whether the recognition result contained any errors (i.e. by selecting yes or no after hearing the audio). As shown in Figure 2(a), the proportion of correct error detection was higher in UncertainSlow (85%) compared to AllNormal (80%), AllSlow (76%), and UncertainBeep (79%). A non-parametric ART test revealed a significant difference ( = 4.48, = 0.087, = 0.005). Post-hoc pairwise comparisons with Bonferroni correction found a significant difference between the UncertainSlow and AllSlow conditions ( = 0.002). In contrast to the previous study by Hong and Findlater (Hong and Findlater, 2018) that reported improved error detection with a slow speech rate, our study did not find a significant difference in error detection between other pairs. However, our results suggest that slowing down the audio playback only when necessary might help users to better detect the presence of errors compared to uniformly slowing down the playback. As shown in Figure 2(a), the variance in per-participant error detection performance was smaller in the UncertainSlow (SD = 11.5) condition compared to AllNormal (SD = 12.9), AllSlow (SD = 15.8), and UncertainBeep (SD = 16.6). This may indicate the confidence score based slowing helped avoid some users avoid substantially lower accuracy compared to the average.
@@
Unfortunately, we lacked sufficient data to reliably analyze the impact of different audio annotations on participants’ ability to identify the specific locations of the errors. This was due to not every participant experiencing a sufficient number of recognition errors in each condition. However, we were did conduct some analysis by aggregating errors across all conditions. We found users correctly located 49% of all errors. Broken down by type of recognition error, they located 2% of insertion errors, 49% of substitution errors, and 62% of deletion errors. Actual substitution and deletion errors were identified by aligning the reference and recognition transcripts using the Levenshtein distance algorithm (Levenshtein et al., 1966). We determined actual insertion errors by manual review. Across all conditions, the ratio of locating errors was 48% for challenging phrases and 52% for common phrases. This indicates participants missed nearly half of the errors in the transcribed text and this was not strongly influenced by the difficultly of the sentence. In particular, participants struggled to identify insertions, suggesting that detecting and correcting added words may necessitate greater attention.
In our study, participants were presented with both common and challenging phrases. The WER was significantly higher at 17% for challenging phrases compared to 12% for common phrases ( = -0.97, ¡ 0.001). This suggests that our approach of using challenging phrases to elicit more recognition errors was effective.
For common phrases, the proportion of correct error detection was higher in UncertainSlow (90.9%) compared to AllNormal (84.7%), AllSlow (83.5%), and UncertainBeep (82.5%). A non-parametric ART test revealed a significant difference (, , ). Post-hoc pairwise comparisons with Bonferroni correction found significant differences between AllSlow and UncertainSlow (), and between AllSlow and AllNormal (). No significant differences were observed between UncertainSlow and UncertainBeep () or between AllNormal and UncertainBeep ().
For challenging phrases, the proportion of correct error detection was also higher in UncertainSlow (79.9%) compared to UncertainBeep (75.9%), AllNormal (75.6%), and AllSlow (68.5%). An ART test revealed a significant difference (, , ). Post-hoc pairwise comparisons with Bonferroni correction found significant differences between AllNormal and AllSlow (), and between AllNormal and UncertainBeep (). No significant differences were observed between UncertainSlow and UncertainBeep (), or between UncertainSlow and AllNormal ().
Comparing results between the common and challenging phrases, it is evident that error detection accuracy decreases for all conditions when the phrases become more challenging. However, the relative advantages of UncertainSlow remained consistent and had the highest detection accuracy across both common (90.9%) and challenging (79.9%) phrases. This suggests that selectively slowing playback in UncertainSlow effectively supports error detection across varying task difficulties.
@@ 3.2. Detection Time
We measured the detection time from the end of the audio playback to the participants’ response of either yes or no. Average detection times were similar: 1.98 seconds in AllNormal, 2.03 seconds in All Slow, 1.86 seconds in Uncertain Slow, and 2.26 seconds in Uncertain Beep condition (Figure 2(b)). These differences were not significant ( = 0.69, = 0.014, = 0.56). However, when we considered the total time to review errors including the audio playback time, we observed a significant difference ( = 14.2, = 0.23, ¡ 0.001). Participants took less time in reviewing errors in AllNormal (4.52 s) condition compared to UncertainSlow (4.85 s), AllSlow (5.43 s), and UncertainBeep (5.42 s) conditions. Post-hoc pairwise comparisons revealed significant differences between AllSlow and UncertainSlow ( ¡ 0.001) but not between any other pairs. This indicates that selectively slowing down the audio did not significantly increase the overall review time despite the increased audio length.
3.3. Subjective Feedback
Participants rated how easy it was to identify errors under four conditions on a 7-point Likert scale (1 = strongly disagree, 7 = strongly agree). As shown in Figure 3, 92% of participants found AllSlow easy, followed closely by 94% for AllNormal, 92% for UncertainBeep, and 86% for UncertainSlow. The Friedman test indicated no significant difference in participants’ ratings across the four conditions ( = 2.94, = 0.40), suggesting that participants did not perceive substantial differences in how easy it was to identify errors under the different conditions.
A horizontal stacked bar plots showing participants’ feedback on how easy it was to identify errors under different conditions. The percentages on the left represent the portion of participants who strongly disagreed or disagreed or slightly disagreed, the percentages in the middle correspond to neutral and the percentages on the right correspond to those who strongly agreed or agreed or slightly agreed with the statement.
@@ 4. Discussion
In our study, we used four audio annotations to assess participants’ ability to detect errors in their transcribed speech. Our result suggests that using the recognizer’s confidence in its results can be used to change how we present the result audio to help user detect errors. One limitation of our study is that we only considered sighted native English users. Blind users, for example, may have more experience listening to TTS, which could impact their ability to detect errors in TTS audio. Additionally, non-native users with ascents may have different experiences with speech recognition technology. Future research should explore how diverse populations detect errors when using speech recognition technology.
We do not know the environment our crowdsourced participants completed our study, but it is likely many were in a quiet environment. In real world use, speech recognition users may be exposed to various types of noises and distractions that could affect their ability to detect errors. Moreover, users may be engaged in other tasks while using voice assistants (e.g. driving or exercising), which could also affect their ability to detect errors. Future studies could investigate how different contexts and tasks impact users’ ability to detect speech recognition errors.
To create a realistic task, we had participants record themselves speaking sentences and used a state-of-the-art speech recognizer to present their actual transcription results. While we could have artificially forced recognition errors, we wanted to study the impact of our interventions in a more externally valid task with accurate speech recognition, but with some phrases containing difficult words (as might happen in real-world use). However, this resulted in our data lacking sufficient recognition errors for each participant and condition for fine-grained analysis of users’ ability to locate errors. We suggest future work consider additional ways to ensure sufficient errors in each condition such as: 1) a longer or multi-session study with more utterances per condition, 2) adding noise to participant’s audio to increase errors, or 3) occasionally presenting the second best recognition result.
@@ 5. Conclusion
Our study investigated whether audio annotations based on confidence scores helped participants detect speech recognition errors. Annotating audio based on confidence scores showed a trend toward improved error detection including reducing the variability in the detection accuracy of our users. However, the 85% detection accuracy of selectively slowing playback was not statistically different than the 80% accuracy of the baseline condition that played all results at normal speed. This highlights the need for further research to validate and expand upon these findings. Nevertheless, we believe our results can help inform the design of more effective error detection features that improve eyes-free speech interaction.